# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ 2.1.3 Human behavioural patterns identification
Here, sigma is a scaling parameter that determines the closeness of data points to be considered a cluster. The other parameter $\mu$ is a predefined value. In this study, it will be referred to as self-similarity, which will be in the range of [0,1].
The Gaussian kernel was selected because it will assign high similarity to close-by points, allowing them to cluster together and assign zero similarity to points far apart. This means that the number of clusters will depend on the sigma value, which is evident from the results.
Figure 6: Sigma sweep for bank worker occupation class.
First, a similarity matrix (Adjacency matrix) was formulated as mentioned below.
@@
    Largest Gap Value: The eigen gap with the highest value was considered as the prominent mode.
@@
    Selection of Dominant Sigma: Once the prominent mode is determined, the sigma value corresponding to the highest gap value for this mode is selected as the dominant sigma according to the selection algorithm [35].
@@
    Clustering Procedure: With the prominent mode and dominant sigma identified, clustering can be executed by following the steps described in [35].
When considering the sigma sweep for the bank worker class, as shown in Fig. 6, the prominent mode could be determined as 23, which would mean the behavior would be broken down into two significant clusters at the corresponding sigma value. Then the matrix,
$X=[v_{1},v_{2},v_{3},...,v_{n}]\in\mathbb{R}^{n\times{n}},$ (5)
can be formulated by stacking eigenvectors in columns. By treating each row of $X$ as a point, the resulting points can be clustered into the desired number of groups using K-means clustering.
@@
2.1.4 Bimodality probability matrix generation
To model human motion, we employed a robust probabilistic approach capable of capturing complex movement dynamics, as introduced by Weligampola et al. [13]. This method enabled the generation of daily trajectories for agents within a predefined environment, taking into account the type of day (weekday or weekend) and the agent’s profession.
The inspiration for adopting this method stemmed from observations of GPS trajectories. As previously observed, denser points in GPS data were the locations where people tended to stay for a longer time. It was further observed that human motion consists of pseudo-random visits to locations and a stay time associated with each location. These observations form the basis of the theory proposed by Weligampola et al. [13].
To model human movement, the theory introduces two probability matrices with size $n\times m$, where $n$ represents the number of zones or locations and $m$ represents the total number of minutes in a day:
@@ 2.1.4 Bimodality probability matrix generation
The $\mathcal{L}$ matrix is influenced by the time of day, while the $\mathcal{S}$ matrix depends on the location being visited. Fig. 7 illustrates how these matrices are utilized to model human motion. At any given moment, an individual is assigned probabilities for being at various locations. The higher the probability, the greater the likelihood of being at that location. The actual location a person visits is determined using the roulette wheel selection method applied to the matrix $\mathcal{L}$. Once a location is selected, the duration of stay at that location is determined using matrix $\mathcal{S}$ in a similar manner. After the stay duration ends, the process repeats with a new location and corresponding stay duration. This iterative approach is used to generate schedules for agents in the second stage of AVSim: ABM simulation.
Figure 7: Process of generating daily trajectories.
Figure 8: Probability variation for the most visited locations with time for Students.
Using the time-location dataset generated earlier, we extracted these two matrices for each profession class. Two rows extracted from a generated matrix $\mathcal{L}$ are illustrated in Fig. 8. This figure shows the visit probabilities for two of the most visited locations for students throughout a regular school day.
2.2 Software system
AVSim was implemented using Python for its flexibility and extensive libraries for simulation and data processing. The class and object structure based on object-oriented principles were used to represent the system’s agents, locations, and transport mediums. This approach improved code modularity, reusability, and ease of maintenance. For data visualization, a React application was employed, utilizing the React-D3-Tree library for visualizing tree structures. Python’s Folium library was used to generate interactive maps.
2.3 Environment
To represent the geographical hierarchy identified in Section 2.1.2, we used a tree-based hierarchical structure to represent the environment within the ABM.
Figure 9: Map illustrating the Kandy District with annotated cities, sub-cities, and functional zones.
This structure is advantageous as it allows agents to navigate between locations in a way that mirrors real-world movement. When an agent transitions from one location to another, the tree structure inherently defines the sequence of locations or zones it must traverse, which is essential for understanding and predicting the spread of disease.
Given that the motion data was sourced from the Kandy District in Sri Lanka, the same environment was recreated within the ABM framework. The district’s geography, including its three major cities—Kandy, Pallekele, and Gampola—was annotated using polygon shapes on Google Maps. Each city was further subdivided into sub-cities and zones, classified based on their primary functions, such as educational, medical, and residential areas. For instance, schools located near each other within a city were grouped into a single educational zone. Fig. 9 illustrates the annotated cities and zones.
The annotations were exported, and the polygon coordinate points were used to construct the environment within the ABM. The hierarchical structure was built using the NetworkX library.
At the base level of the hierarchy, the tree’s leaves represent primary locations such as schools, homes, and hospitals. These locations were randomly placed within designated zones (e.g., educational, residential, medical) to reduce manual annotation, as specific placements within zones have minimal impact on disease spread. The zones form the next hierarchical level. To differentiate locations and zones, a numbering system (e.g., "Home_1," "Home_2," "ResidentialZone_1," "ResidentialZone_2") was applied. These names were derived from the generated dataset to align with the created bimodal probability matrices.
Higher levels of the hierarchy represent broader geographical categories such as sub-cities, cities, and districts. ABM users can define areas based on their needs, regardless of the generated dataset. Fig. 10 illustrates the tree structure, with only a single zone expanded to show its locations for clarity.
@@ 2.3 Environment
In this tree structure, each node represents a geographical area. Nodes are implemented as objects, with attributes defining each area’s characteristics, including its boundary (polygon coordinates), centroid, and the list of agents present. Each node also contains a dictionary to track the number of agents and their respective disease states.
2.4 Vector patch
In agent-based models of vector-borne diseases, it is not essential to represent each vector individually, as is done for human agents. Instead, modeling vector density within localized areas provides a simpler yet accurate representation of the vector population. Following approaches from similar studies [14], the simulation adapts a patch-based approach to represent vector densities.
Each zone is divided into patches, with the patch size determined by the approximate habitat areas of the vectors. Fig. 11 presents a map of Kandy City, illustrating an example of vector density across different patches.
A patch is modeled as an object that encapsulates environmental factors such as temperature, humidity, and rainfall, which influence vector emergence and death rates. It also maintains the number of vectors in different disease states, a dictionary of agents with their disease statuses, and a list of locations within the patch.
Figure 11: Map of Kandy District showing the zones and their patches, with a color bar representing an example vector density across the patches
The total birth rate of vectors in a given patch $k$, $h_{v}^{k}$, is calculated as follows:
@@ 2.4 Vector patch
  • $\psi_{v}$: Average natural emergence rate of vectors
  • $\mu_{v}$: Average death rate of vectors
  • $K_{v}$: Carrying capacity of the vectors in the patch
  • $N_{v}$: Total number of vectors in the patch
2.5 Agent generation and emulation
The agents of the system are a vital component in any agent-based model. In this model, agents were initially categorized based on their occupation class and further subdivided as per the identified mobility patterns. For instance, doctors who worked the day shift vs. the night shift were classified into two groups. Each occupation class was defined as either an adult or child class. This parameter was important to establish the independence of each occupation class. The occupation classes, sub-classes and composition of agents to be modeled by the ABM could be varied based on input data.
The agents were represented through the use of Object Oriented Programming (OOP) concepts, specifically classes and objects. The agent class represented the structure of all agents (objects) created for emulation. Parameters of an agent defined the characteristics of each individual agent, such as the occupation class, sub-class (if applicable), age ranges, home and work location classes, and factors necessary for disease modeling. The total count of agents was dependent on the carrying capacity of the environment. As mentioned in the above sections, this study was capable of modeling the Kandy district, and approximately 2000 agents were generated for emulation. This constituted about 1% of the actual population of the district. The composition of agent occupation classes was also selected to approximately represent their real-life counterparts.
The procedure of agent generation within the ABM was as follows. At the start of emulation, the ABM was given the carrying capacity and composition of agents from the environment and input data, respectively. The agent objects were then generated with parameters set within defined ranges for the respective occupational classes. Subsequently, each agent object was assigned a home object. The occupancy limit for home objects is defined by the emulation parameters. Each house object was constrained to either be empty or to contain one or more adult agents. This was due to child agents not being considered independent in the ABM. Finally, a work location object corresponding to the agent’s work location class was assigned to the agent at random.
Upon completion of the agent generation procedure, the emulation process was initiated. Simulation steps were executed every 1 minute for each day (resolution of the simulation. At the start of an emulation day, a timetable was generated for each individual agent (this process is described in Section 2.1.4). This timetable defined the routine (in terms of locations to visit and time of occurrence) for each agent. Except for home and work locations, all other locations (such as shopping malls and banks) were given as location classes. Therefore, the agent would choose to visit the closest location instance of that class. Once the timetables were generated, each agent would utilize modes of transport (as explained in the next section) to follow their schedule.
2.6 Transport
In the context of disease modeling, long-range human mobility plays a critical role in facilitating the rapid geographical dissemination of emergent infectious diseases [36]. This study incorporates the simulation of long-range human mobility by modeling both private and public transportation systems to better understand their impact on disease spread.
In the private transport modeling approach, a subset of agents was designated with a public transport medium flag, while others were assigned private transport objects at the simulation’s outset. These private transport objects enabled agents to travel to their destinations whenever needed. Conversely, agents marked with the public transport medium flag utilized public transport for all instances of long-range mobility. The primary objective was to evaluate the impact of private transport users on disease propagation.
Public transport modeling holds significant importance in the context of disease modeling, particularly due to the increased vulnerability of passengers to airborne disease transmission in confined spaces with close proximity to others [37]. Agents not assigned a private transport medium flag relied on public transport for long-range mobility instances.
In this study, buses and taxis were used as the primary modes of public transport. Two types of bus objects were employed to represent different hierarchical levels in the simulated environment. Inter-city buses facilitate transport between cities, while intra-city buses facilitate transport inside a city. This process is explained in Fig  12.
Bus objects were initiated at the start of the simulation. Each bus station referred to as a Bus Node, maintained an updated dictionary of available buses and their next destinations.
Figure 12: Public transport mechanism.
When a bus arrived, agents who had been waiting for transportation were assigned to that bus, provided the bus’s next destination matched the agent’s next location in the bus’s dictionary of destinations.
If the agent had to wait too long for a bus, they would be assigned to a taxi to travel to the next destination. The Taxi was assigned to each zone so that every agent who missed the bus could still get a taxi. When a taxi arrived at a destination zone, it stayed there until the next trip was assigned.
2.7 Markov disease progression model
A person can transition through various states after being exposed to a disease. These transitions depend on the nature of the disease. Generally, an exposed individual becomes infectious after a certain period, carrying the risk of spreading the disease. The disease progression may remain asymptomatic or exhibit symptoms, potentially progressing through severity stages such as mild, severe, and critical.
We used a Markov model to handle state transitions, as illustrated in Fig. 13, which defines the possible states a person can be in after exposure to a virus. Transitions between these states were probabilistic and were often modeled using distributions such as the log-normal distribution. The probabilities of these transitions depended on factors like the nature of the disease, age, prior infections, and the quality of medical care. For our two use cases, dengue and COVID-19, the transition times are shown in the figure and followed a log-normal distribution. Values for COVID-19 were taken from Covasim [7] while dengue was taken from WHO [38].
It is important to note that the transition from the "susceptible" to "exposed" state was not considered part of the Markov model. This transition depends on how a person becomes exposed to the virus. In airborne diseases, this typically occurs through human-to-human contact, while in vector-borne diseases, it happens via interactions with vectors. These mechanisms will be detailed in subsequent sections.
@@ 2.7 Markov disease progression model
In this model, we had $N$ possible states (without "susceptible"), and transitions between these states were represented by the transition matrix $A$, where $A_{ij}$ denotes the probability of transitioning from state $j$ to state $i$. We used a probability vector $p_{t}$ (of size $N\times 1$) that captures the state probabilities at time $t$. To find the next state probabilities $p_{t+1}$, we applied the fundamental equation 7 for a Markov chain.
$p_{t+1}=A\cdot p_{t}$ (7)
To determine the actual next state, we generated a random number $r$ between 0 and 1. Using $p_{t+1}$, we calculated cumulative probability ranges for each state $i$, based on the probabilities in $p_{t+1}$. The next state is chosen as the one for which $r$ falls within the cumulative range associated with that state.
2.8 Air borne disease modeling
2.8.1 Disease progression mechanism
Initially, a small percentage of the population was infected with the COVID-19 virus, and the agents were freely allowed to go about their daily business, allowing us to study the disease spread. Once an agent is infected, it will progress through different stages of the disease. This was modeled using the Markov model as shown in Fig  13. Each agent will belong to either susceptible, exposed, infectious, symptomatic, asymptomatic, or dead states.
The disease is spread when a susceptible individual comes into contact with an infected one. Contacts occurred in two scenarios: at specific locations or during transport (on a bus). At each time step, a 1-meter radius around each agent was monitored. If another agent enters this radius, it is recorded as a contact. When a susceptible agent contacts an infected agent, a random number is generated. If this number recedes the agent’s infectious probability ($\rho$), the susceptible agent’s disease state changes to "exposed." Each agent’s disease state is then updated according to the disease model. Agent’s infectious probability depends on the individual’s immunity level, as described in the following section.
2.8.2 Immunity modeling
Contact between a susceptible agent and an infected agent does not always result in disease transmission. This is determined by a transmission probability, which depends on the agent’s immunity. In the simulation, agents were assigned varying immunity levels based on factors such as age, hygiene practices (e.g., maintaining a safe distance, hand washing, sanitizer use), and vaccination status. This infection probability, $\rho$, can be modeled using the equation 8.
@@ 2.8.2 Immunity modeling
where $S_{age}$ is the susceptibility based on the age, $k$ is a tunable parameter that has to be manually found, $\alpha_{vacc}$ is the Weight assigned to vaccination, $\gamma_{vacc}$ is the Immunity boost provided by vaccination, $\alpha_{hyg}$ is the Weight assigned to hygiene factor and $\gamma_{hyg}$ is the Immunity boost provided by safe hygiene practices which include washing hands, wearing masks, and social distancing. The age-based susceptibility values were taken from Covasim [7], and they are presented in Table  2. Through some prior tuning process, the value of $k$ was selected as 0.3.
Table 2: Age-based susceptibility values for agents
Age 0-9 10-19 20-69 70-79 80+
Susceptibility 0.34 0.67 1.00 1.24 1.47
2.8.3 PCR testing and quarantine mechanism
@@
Polymerase Chain Reaction (PCR) testing is crucial in predicting the risk of airborne diseases [39]. This is used as a pre-event testing activity, and people with positive PCR results are identified as infected agents. In the COVID-19 pandemic, infected agents were identified using PCR tests and using contact tracing methods. Contacted agents were subjected to quarantine or self-isolation [40].
The AVSim had the functionality to perform PCR tests to identify the infected agents. Each agent would have a daily testing probability, which would depend on the state of infection. This method allowed the user to specify the testing probabilities for each of these states. These numbers would depend on the regulations of the government and the healthcare sector. Therefore, the simulation would decide the number of PCR tests done each day.
Two methods were employed in the simulation to perform quarantine. In the first method, if a PCR test was positive, the relevant agent would be isolated and put under quarantine. When that agent was recovered and safe to interact with other agents, the simulation would allow the agent to end their quarantine. In the second method, all agents within a specific occupation class were quarantined if the number of infected agents in that class surpassed a certain threshold. In 3.3, results for each of these methods are analyzed and discussed.
2.8.4 Vaccination mechanism
Vaccination mechanisms are utilized to enhance community immunity, particularly during outbreaks of airborne diseases like COVID-19 [41]. In AVSim, vaccination events were organized to reduce the disease spread while increasing the immunity levels of the agents.
@@ 2.8.4 Vaccination mechanism
In AVSim, an outbreak was declared whenever the total number of infected agents exceeded a specified threshold. When an outbreak was detected in the simulation, vaccination events could be organized. The AVSim simulation allowed the organization of vaccination events on a selected day at a specified geographical zone. AVSim further allowed the capability to vaccinate targeted occupational classes and gave out various types of vaccines. When an agent is given a vaccination, their vaccine immunity $\gamma_{vacc}$ is boosted. According to  [42], a few possible vaccine types and their immunity boosts are shown in Table  3. New vaccination events with new vaccine types could be added this way. Additionally, vaccination could be administered in doses.
Table 3: Different types of possible vaccines and their effectiveness
Vaccine Name Immunity boost after 1st dose Immunity boost after 2nd dose
Moderna 74 93
AstraZeneca 78 67
Pfizer 84 93
2.9 Vector-borne disease modeling
@@
For modeling vector-borne diseases, the network-patch approach proposed by Carrie Manorea et al. [14] for dengue transmission has been adapted in this study to reflect the dynamics of vector-borne diseases more accurately.
2.9.1 Disease progression in vectors

### v2

@@ III-A2 Behavioral patterns clustering
Here, sigma ($\sigma$) is a scaling parameter that determines the closeness of data points to be considered a cluster. The other parameter, $\mu$, is a predefined value. In this study, it will be referred to as self-similarity, which will be in the range of [0, 1]. In our case, we selected $\mu$ as 1 since it yielded the best possible results.
First, a similarity matrix (Adjacency matrix) was formulated as shown below.
@@
Figure 4: Sigma sweep for bank worker occupation class.
@@
    The eigen gap with the highest value was considered as the prominent mode.
@@
    Selection of Dominant Sigma: Once the prominent mode is determined, the $\sigma$ value corresponding to the highest gap value for this mode is selected as the dominant $\sigma$ according to the selection algorithm [31].
@@
    Clustering Procedure: With the prominent mode and dominant $\sigma$ identified, clustering can be executed by following the steps.
For example, when considering the sigma sweep for the bank worker class, as shown in Fig. 4, the prominent mode could be determined as 2-3, which would mean the behavior would be broken down into two significant clusters at the corresponding $\sigma$ value. Then the matrix,
$V=[v_{1},v_{2},v_{3},...,v_{n}]\in\mathbb{R}^{n\times{n}},$ (5)
can be formulated by stacking eigenvectors in columns. By treating each row of $V$ as a point, the resulting points can be clustered into the desired number of groups using K-means clustering.
@@
III-A3 Modeling agent trajectories
To model human motion trajectories, we employed a robust probabilistic approach capable of capturing complex movement dynamics. This was introduced as a hypothetical model for modeling human motion by Weligampola et al. [3]. We enhanced and validated it using real-world GPS data, allowing us to generate daily agent trajectories based on day type (weekday or weekend) and profession within a defined environment.
Figure 5: Process of generating daily trajectories.
This method was inspired by patterns observed in GPS trajectories, where denser point clusters indicated longer stays. It was also noted that human movement involves pseudo-random visits to locations with varying stay times. These observations form the basis for the theory proposed by Weligampola et al. [3].
We start by introducing two probability matrices with size $n\times m$, where $n$ represents the number of locations and $m$ represents the total number of minutes in a day:
@@ III-A3 Modeling agent trajectories
We extracted these two matrices for each profession class using the time-location dataset generated earlier. As shown in Fig. 5, we generate trajectories through roulette wheel selection on these matrices.
III-B Environment
Given that the motion data was sourced from the Kandy District in Sri Lanka, the same environment was recreated within the ABM framework. Major cities such as Kandy, Pallekele, and Gampola were modeled using polygons on a real-world Map and further divided into sub-cities and functional zones (e.g., educational, medical, residential). For example, nearby schools were grouped into one educational zone, as shown in Fig. 1 (c).
To represent this geographic hierarchy, we used a tree-based structure in the ABM, enabling agents to move realistically between locations. Each node in the tree represents an area and is implemented as an object with attributes defining its characteristics (Fig. 6, showing one expanded zone for clarity).
@@ III-B Environment
III-C Agent generation and simulation
In AVSim, agents were first grouped by occupation and then further divided based on behavioral patterns. For example, day-shift and night-shift doctors were treated as separate groups. Each agent was then assigned parameters that defined their characteristics, including occupation class, sub-class (if applicable), age ranges, home and work location classes, and other factors necessary for disease modeling. This agent composition was selected to approximately represent their real-life counterparts. Thereafter, the simulation process was initiated. Simulation steps were executed every 1 minute for each day (resolution of the simulation). Each day began with timetable generation for agents (see Section III-A3), after which they followed their schedules using available transport modes.
III-D Transport
In the context of disease modeling, long-range human mobility plays a critical role in facilitating the rapid geographical dissemination of emergent infectious diseases [32]. Our study incorporates the simulation of long-range human mobility by modeling both private and public transportation systems to better understand their impact on disease spread.
Whenever an agent had to travel from one location to another, the traveling path in the simulation environment(intermediate cities and zones the agent passed by) was obtained using the graph object shown in Fig  6. For this, the ”traveling salesman problem” algorithm  [33] was used.
Public transport modeling holds significant importance in the context of disease modeling, particularly due to the increased vulnerability of passengers to airborne disease transmission in confined spaces with close proximity to others [34].
In this study, buses and taxis were used as the primary modes of public transport. Two types of bus objects were employed to represent different hierarchical levels in the simulated environment. Inter-city buses facilitate transport between cities, while intra-city buses facilitate transport within a city. This process is explained in Fig. 7.
Figure 7: Public transport mechanism.
III-E Vector patch
In agent-based models of vector-borne diseases, it’s unnecessary to model each vector individually. Instead, representing vector density within localized areas offers a simpler yet effective approach. Aedes aegypti mosquitoes, for example, usually stay near human habitats and travel up to 400 meters to lay eggs [35]. Based on similar studies [8], we use a patch-based method to model vector density.
Each zone is divided into patches, sized according to typical mosquito habitat ranges as shown in Fig. 1 (d).
Each patch is characterized by environmental factors such as temperature, humidity, and rainfall, which influence vector emergence and death rates. The total birth rate of vectors in a given patch $k$, $h_{v}^{k}$, is calculated as follows:
@@ III-E Vector patch
Here, $\psi_{v}$ is the average natural emergence rate, $\mu_{v}$ is the average death rate, $K_{v}$ is the carrying capacity, and $N_{v}^{k}$ is the total number of vectors in patch $k$.
III-F Markov disease progression model
After exposure to a disease, a person transitions through different states based on factors such as disease type, age, prior infections, and medical care. Typically, an exposed individual becomes infectious, potentially spreading the disease asymptomatically or with symptoms ranging from mild to critical. We can model this using a Markov model (Fig. 8), with transition times following a log-normal distribution [3]. Data for COVID-19 were taken from Covasim [6] while dengue was taken from WHO [35].
@@ III-F Markov disease progression model
III-G Airborne disease modeling
III-G1 Disease progression mechanism
Initially, a small percentage of the population was infected with the COVID-19 virus, simulating natural disease spread as they moved freely. Transmission occurred either at specific locations or on public transport. At each time step, a 1-meter radius around each agent was monitored. If another agent enters this radius, it is recorded as a contact. Disease transmission between agents is governed by their transmission probabilities, $\rho$, which is influenced by the agent’s immunity. In the simulation, agents were assigned varying immunity levels based on their age, hygiene practices (such as social distancing, hand washing, and sanitizer use), and vaccination status. The infection probability is modeled as:
@@ III-G1 Disease progression mechanism
where $S_{age}$ represents age-based susceptibility, $\alpha_{vacc}$ and $\alpha_{hyg}$ are the respective weights for vaccination and hygiene, and $\gamma_{vacc}$ and $\gamma_{hyg}$ represent the immunity boosts provided by vaccination and hygiene practices. $k$ is a tunable parameter.
Values for $S_{age}$ were taken from Covasim [6]. Through some prior tuning processes, the value of $k$ was selected as 0.3.
III-G2 PCR testing and quarantine mechanism
@@
Polymerase Chain Reaction (PCR) testing is crucial in predicting the risk of airborne diseases [36]. In AVSim, this could be used as a pre-event testing activity, and people with positive PCR results were identified as infected agents. In the COVID-19 pandemic, infected agents were identified using PCR tests, and exposed agents were identified by tracing the paths of contacts of infected agents. Both infected and exposed agents were then subjected to quarantine or self-isolation. Therefore, AVSim could be used by policymakers to study the effect of various PCR and quarantine policies.
III-G3 Vaccination mechanism
@@
Vaccination mechanisms are utilized to enhance community immunity, particularly during outbreaks of airborne diseases. In AVSim, vaccination events could be organized to reduce the disease spread while increasing the immunity levels of the agents. The AVSim simulation allowed the organization of vaccination events on a selected day at a specified geographical zone. AVSim further allowed the capability to vaccinate targeted occupational classes and give out various types of vaccines. When an agent is given a vaccination, their vaccine immunity $\gamma_{vacc}$ is boosted.
III-H Vector-borne disease modeling
@@
For modeling vector-borne diseases, the network-patch approach proposed by Carrie Manorea et al. [8] for dengue transmission has been adapted in this study to reflect the dynamics more accurately. The transport model was excluded, as the likelihood of contracting a vector-borne disease during transit is minimal. Nevertheless, the agents’ motion dynamics were retained.
III-H1 Disease progression in vectors